In February the Anadromous Program continued winter work mainly consisting of policy and technical work. Participating in several policy processes including the Spokane River Lead Entity development, Columbia Basin Collaborative, Blocked Areas Anadromous Fish Working Group, and the Columbia River Treaty. Staff continued bi-weekly monitoring of juvenile chinook in net pens in Lake Roosevelt. Prep work for spring monitoring and construction of equipment is ongoing. Installation of monitoring equipment has begun.
The month of February has been busy with fisheries focused on equipment preparation and maintenance, continued fish health assessments of net pen fish. Safety and equipment training is ongoing for new staff and those in need of renewals. We continue our participation in regional and national processes and workgroups. Several grant proposals have been awarded and contracts are being finalized. Staff attended aquaculture meeting and prepared grant and contract budgets.
Spokane Tribal Hatchery Program
Fish production consisting of:
- 13,800 rainbow trout adults
- 40,000 rainbow trout juveniles
- 985,000 rainbow trout fry
- Monthly EPA National Pollution Discharge Monitoring Report
- Revised 2024 BPA Work Element Budget for increased funding.
- Assisted snow plowing, ongoing operation and maintenance of fish hatchery facilities and grounds.
Spokane Tribal Park Rangers Program
The Ranger department and Some DNR staff have actively been working on and are near completion of suggested revisions to Ch 17 of the Spokane tribal law and order code. Ranger department participated in the annual lake Roosevelt ARPA webinar hosted by the national park service. The cold snap has slowed the amount of angler checks for the month. The Bank Fishing has been good and will resume with warmer weather.
The drawdown has begun, and rangers are monitoring the beaches for ARPA violations. Wood cutters are still active in the cottonwood area and Wellpinit mtn logging unit. Predator hunters been in pursuit of Coyotes and cats after every fresh snow. Shed antler season has begun and will continue through the spring months.
All rangers are scheduled to attend the TERO sponsored CPR/AED/First aid training next month. A new patrol vehicle should be in service near the Middle of march. The Duckworth jet boat is in for repairs, The other boat will be placed in service soon after the marina is ice free.
Wildlife and Timber, Fish and Wildlife Programs
In February the Wildlife Program has been primarily focused on trying to capture mule deer. That included setting and checking clover traps and searching for mule deer to free range dart. There are 7 mule deer collars to deploy and so far, we have not been successful. We have caught white-tailed deer and elk in our traps. We are also prepping for the upcoming public wildlife committee meeting at the beginning of March. This includes updating a power point presentation the Wildlife Program gives, with all hunt data, aerial surveys, and project updates. Other office work has included budget modifications, ordering supplies, and entering habitat survey data.
During the month of February, the program has completed building Clover Traps, the program has deployed clover traps, monitoring has been ongoing throughout the month. Staff have been involved in various trainings throughout the month. Shop projects are ongoing. Staff have had two meetings with BPA to discuss contract execution and Land Swap Efforts.
The timber, fish, and wildlife program continued work to monitor timber harvest activities on private and state lands within the Tribe’s Usual & Accustomed area. This included review of 10 Forest Practices Applications in Stevens, Spokane, and Lincoln counties. This work helped to achieve adequate natural resource protections and compliance with state regulations. Additionally, the program continued its collaboration with the Washington Forest Practices Adaptive Management Program (AMP) through participation in the Scientific Advisory Group Eastside (SAGE), the Instream Scientific Advisory Group (ISAG), the Cooperative Monitoring, Evaluation, and Research Committee (CMER), and the Washington State Forest Practices Board. Participation involved representation in meetings as well as review of AMP documents, including the Water Typing Strategy Charter. The CMER Water Typing Strategy includes study development for Potential Habitat Breaks (PHB), Default Physical Criteria (DPC) assessment, LiDAR based water typing models, and fish/habitat detection using eDNA.
The Range program continued entering data from various vegetation monitoring plots across the reservation, organized GPS data, and worked on data analyses. This all will help guide projects for the 2023 field season. Local meetings were attended regarding noxious weed control, where the Range program developed several regional partnerships within the Columbia River Basin. The program continued working on projects planning for the upcoming field season, including tree and shrub planting, grass planting, and weed control. The program, with assistance from Federal partners, conducted a feral horse survey across the reservation.
Air Quality Program
- Radon Poster still running till the end of March.
- Travel to Washington, DC for National Tribal Air Association.
- National Tribal Forum on Air Quality-Helping with set up for 2023.
- 2/6- Internal staff meeting- Planning human detections dogs, collections need, river projects.
- 2/8- BPA, BOR quarterly meeting all day zoom.
- 2/10- MAC land acknowledgment with Gena Peone & Tisa Matheson- zoom meeting.
- 2/14- Lake Roosevelt Project Specific Programmatic agreement
- 2/16- Zoom meeting with BOR, BPA, NPS- Traditional cultural property meeting.
- 2/20- Human detection dog bid to contactors
- 2/23- MAC land acknowledgment with Gena Peone & Tisa Matheson- zoom meeting.
- HVAC meeting with McKinstry
- 2/27- ARPA meeting with NPS, STI, CCT-
- Wake boat meeting with BOR- STI- CCT- NPS
Collections staff assessed student engagement during school presentations, along with the BPA/BOR Deliverable for public education and outreach, and subsequently updated the STIPP Collection and Archives PowerPoint presentation to realign it with STIPP’s outreach program goals. The updated PowerPoint has been used twice this month during presentations with third-grade students at the Libby Center and Roosevelt Elementary School with noticeable improvement in student engagement and interest. Collections staff has also committed time to review and plan building upgrades with a focus on safety for staff and collection materials. Upgrades that are currently in the planning and scheduling phase include inspection of fire extinguishers, electrical updates, a building inspection, and an HVAC upgrade to address humidity control.
- Waiting for comments from local governments on the Notice of Application Letters sent out for the Spokane Tribe’s Fee to Trust 155.33 acres in Airway Heights, WA.
- Working with AVSO appraisers to setup conference call with Gregory Richards (Contractor) on 2 appraisals for the Lagoon Right of Way.
- Working on the Land Titles & Records Project to submit documents and affidavits correcting Title Status Reports.
- Completed 2 land conveyances and requested NEPA Cat Ex for them through NEPA Tracker.
- Submitted 2 home site lease requests to the IRMP ID Team for their review/comments.
- Completed 2 home site leases/approved by resolution-TBC and Acting Superintendent.
- Bureau of Land Management reviewed 5 legal descriptions for us, which will now be used for land conveyances.
- Met with the Acting ED/Planning Director and a landowner regarding a possible land purchase in the Wellpinit Area.
- Met with the Acting Superintendent regarding the Tribe’s Fee to Trust 155.33 acres in Airway Heights and his signature on the numerous Affidavits for the Land Titles & Records Project.
- Working with Stevens County Title Co. on obtaining Commitment of Titles for the Tribes purchase of undivided fee interests.
Forestry and Fire Programs
- Prepared sale documents for tribal and allotted timber sales
- Invoiced for the volume and stumpage for active sales.
- Finalized Cutting Agreement
- Continued development of updated Forest History
- Re-initiated TFMS Scaling Software finalization.
- NWRO Fuels Workshop February 14-16
- NWRO Forest Development Winter Meeting February 14-16
- Full road restriction went into effect, all timber sales temporarily shut down (Breakup). Contractors did get most of their loads to the mills before the shutdown.
- Firebreak sale is close to completion.
- Frost Pocket: The sale made great progress before the closure due to the freezing conditions this past winter. The sale looks to be completed by the expiration date.
- Allotment 2021: One small homesite remains.
- Mima Mound: Sale ready to begin once Break-up is lifted.
- 26 Active Forestry Work Contracts, no progress due to snow depth and limited access.
- Attended Regional Forest Development Meeting 2/14-2/16.
- Upcoming 2 Day Reforestation Conference 3/7-3/8; travel status x3
- Dodge plow truck back from the shop, still under conditional operation to break the rebuilt transmission in.
- GIS Organization continues.
- In process of finalizing a Nursery Services contract with PRT. Working with the PRT and Contracts/Grants to develop a contract that satisfies both parties.
- New tire for George’s work truck due to puncture that cannot be repaired.
- Reformatted FD Technician Job Descriptions, 1 signature pending before being finalized and re-submitted.
- Other daily tasks as needed.
Work continues for the presales crew marking the Crow Hop Project Area.
- Rail Creek prescriptions finished and signed.
- Hog Farm FOR Completed and Signed
- Attended INGY co-op meeting.
- Archiving sales that have been sold recently.
- Started recon on The Docks Project Area
Fire Management Program
- GIS meeting at fire management and working on Acuity international to get the fire physicals for all red card holders.
- working on Acuity International sending out more physicals
- FMO regional meeting
- meeting at Bruce Wynne room
- working on pre work for a training I am going to.
- All thin/chip project work (crew/contracts) under snow closure again
- Budget Mod re-submitted per Compliance request.
- Updating 4 Yr. Plan of Work in NFPORS per BIA guidance
- 1 Staff completed S131 FFT1 training.
- Ongoing vehicle/equipment readiness and repairs
- finished processing orthoimagery into various mosaic scales and resolutions (will be dispersed this week)
- created several requested allotment maps for realty.
- began parcel layer update.
- began researching digitization methods for historic topographic maps of the reservation from the 1930s.
- troubleshot access to Gisd001 buffalo drive for multiple computers.
- continued correspondence with Frontier Precision for GPS units